OpenStudy (anonymous):

What is the simplified form of x^2-25 over x^2-3x-10? A. x-3 over x-5 B.(x-5)(x+5) over (x+3)(x-3) C. x+5 over x+2 D. x-5 over x+2 A medal for the best answer.

OpenStudy (ryan123345):

Ok do you have an idea how to simplify \[x^2-25\]

OpenStudy (anonymous):

None. Im brain dead today.

OpenStudy (ryan123345):

ill give you a hint. What are the two numbers that multiply to 25 and are the same?

OpenStudy (anonymous):

5 right?

OpenStudy (ryan123345):

Good so we set it up as (x 5) (x 5) what do you think the signs would be?

OpenStudy (anonymous):

+ and +?

OpenStudy (ryan123345):

No. - +

OpenStudy (anonymous):

I told you I was brain dead sorry

OpenStudy (ryan123345):

so you get \[\frac{ (x-5)(x+5) }{ x^2-3x-10 }\]so now we got to factor x^2-3x-10

OpenStudy (ryan123345):

what two numbers multiply to 10 but subtract to 3?

OpenStudy (anonymous):

I don't like riddles

OpenStudy (ryan123345):

try it what are the factors of 10? 10,1 2,5

OpenStudy (ryan123345):

which ones subtract to 3?

OpenStudy (anonymous):

2 and 5

OpenStudy (ryan123345):

good. so now you factor like this (x-5)(x+2) and from there you get this \[\frac{ (x-5)(x+5) }{ (x-5)(x+2) }\]which two parentheses cancel out?

OpenStudy (anonymous):

x-5

OpenStudy (ryan123345):

nice job, and then you get?

OpenStudy (anonymous):

x+5 and x-2

OpenStudy (anonymous):

x+2

OpenStudy (ryan123345):

good \[\frac{ x+5 }{ x+2 }\]

OpenStudy (ryan123345):

And of those answer choices the same? :)

OpenStudy (anonymous):

thanks a lot :)

OpenStudy (ryan123345):

No Problem